# Gravitational waves of quasi-circular, inspiraling black hole binaries in an ultralight vector dark-matter environment

### v1

@@ I Introduction
Although there is strong evidence that dark matter represents approximately $25\%$ of the energy content of the universe today, its nature still remains unknown. Many dark-matter models (be it wave-like or particle-like) have been considered, with a wide range of masses and interactions. The lightest class of candidates is usually referred to as ultralight dark-matter field models, where the mass of the field that describes the dark matter ranges from $(10^{-28},1)\;\rm{eV}$. These candidates have gained attention in recent years due to their agreement with the fiducial $\Lambda$CDM model at large scales and their different predictions at small scales [1, 2, 3, 4].
One distinctive feature of ultralight dark-matter candidates is that they induce oscillatory gravitational potentials. The frequency of the oscillations is determined essentially by the mass of the ultralight dark-matter particles. In cosmological contexts, the dark-matter field oscillations usually average out, but this need not be the case around astrophysical systems. In fact, previous work has shown that an oscillatory ultralight, dark-matter environment induces oscillatory behavior in the spacetime metric of astrophysical systems, for example impacting the orbital dynamics and evolution of binary pulsars [5, 6, 7]. Could such an ultralight, dark-matter environment also affect the gravitational waves emitted by compact binaries?
General relativity predicts that gravitational waves are excited by any system with a non-trivial, time-varying multipole-moment structure, such as due to the orbital acceleration of a binary system. When in isolation, the orbital dynamics and evolution of inspiraling compact binaries can be determined from perturbative solutions to the Einstein equations. In practice, when the compact binary is sufficiently well-separated, one employs a post-Minkowskian expansion (in weak fields) or a post-Newtonian expansion (in both weak fields and small velocities) to calculate the gravitational waves in the inspiral phase of coalescence from first principles [8]. When not in isolation, for example due to the presence of an accretion disk [9, 10, 11] or a dark-matter environment [12, 13], the orbital dynamics and evolution of inspiraling compact binaries is in general modified, leaving an environmental imprint on the gravitational waves generated. The question is then not whether an ultralight, dark-matter environmental imprint is present in gravitational waves, but rather whether this imprint is detectable.
@@
Recent efforts have also been made to search for the direct imprint of ultralight dark matter in gravitational waves. For example, in [19, 20], the dephasing of gravitational waves produced by dark-matter fields was considered. In these works, the dephasing is produced in the propagation of the gravitational-wave signal from the source to the detector, caused by the gravitational potentials generated by the ultralight dark-matter fields. In this work, however, we focus on how these gravitational potentials create a dephasing in the generation of the gravitational-wave signal, by changing the world-line of the binary system components, rather than by correcting the way the gravitational waves propagate.
@@
When the binary systems emitting gravitational waves is immersed in a virialized ultralight, dark-matter halo, the dark matter that forms the halo (be it scalar or vector particles) can be described as a collection of waves with random phases [21, 22, 23, 24]. These waves induce oscillations in the spacetime metric, which then perturb the orbital motion of the binary. In particular, the ultralight dark-matter oscillations induce a correction to the acceleration of the binary components, which then affects the orbital energy and angular momentum, introducing dark-matter corrections to Kepler’s third law. The ultralight dark-matter environment, therefore, perturbs the orbital motion of the binary, which then leaves an imprint on the time-varying multipole-moment structure of the spacetime, and thus, on the gravitational waves emitted. Given this, one expects that the strength of the dark-matter imprint on the emitted gravitational waves to be proportional to the dark-matter density and sensitive to the dark-matter mass, which sets the dark-matter oscillation scale, specially close to resonances between this oscillation scale and the orbital timescale.
@@
As we show in the paper, the strength of the ultralight dark-matter effect on the gravitational waveform is proportional to the energy density $\rho_{\rm{DM}}$ of the dark-matter field in the surroundings of the black holes binary. Several works [25, 26] have estimated the dark-matter density in our Solar System to be $\rho_{\rm{DM}}\sim 0.3\rm{GeV}/\rm{cm}^{3}\sim 0.01\rm{M}_{\odot}/\rm{pc}^{3}$. However, the dark-matter density distribution in galaxies is still a matter of debate. In particular, the dark-matter distribution surrounding black holes remains uncertain and varies depending on the dark-matter model. For example, various works [27, 28, 29, 30, 31] suggest that scalar field models spike near black holes, where the density may be several orders of magnitude bigger than the one estimated for our Solar System. We find that the vector ultralight dark-matter corrections are indeed proportional to the dark-matter density, and inversely proportional to a certain power of the ultralight dark-matter mass and the total mass of the binary. Moreover, we find that when the gravitational-wave frequency exceeds a certain threshold, given by a stationary point condition that depends on the vector field mass, the ultralight dark-matter corrections are strongly suppressed. On the other hand, when the gravitational-wave frequency is below this threshold, the vector ultralight dark-matter effect on the gravitational-wave phase for quasi-circular orbits is partially degenerate with the time and phase of coalescence of the binary, rendering this effect hard to measure. This leads to a Heaviside-type modification to the Fourier transform of the gravitational-wave phase, and it is this sudden turn-on/turn-off nature of the modification that yields the dark-matter effect potentially measurable.
Our work suggests that chirping, massive binaries that sample a range of frequencies in the LISA band (and thus, that cross the stationary point condition activating the Heaviside dark-matter modificatons) are necessary to observe this type of dark-matter effect. Through a Fisher analysis, we estimate that LISA would be able to constraint the ultralight dark-matter mass range $(10^{-19},10^{-16}){\rm{eV}}$ for gravitational waves emitted by massive [$(10^{2},10^{6})\rm{M}_{\odot}$] black hole binaries of comparable mass, provided the ultralight dark-matter density is above $(10^{14},10^{15})\;\rm{M}_{\odot}\;{\rm{pc}}^{-3}$. Our work, therefore, suggests that future LISA observations of the gravitational waves emitted in the quasi-circular inspiral of massive black holes may detect or yield interesting constraints on the density of vector-type ultralight dark matter.
@@ II Ultralight dark-matter vector field around a compact binary
In this section, we present the vector-field dark-matter model that we study in this paper, following mostly [7], for completeness and to establish notation. Let us then describe the dark matter through an ultralight vector field minimally coupled to gravity, whose action is
[…]
@@
We consider that the vector dark matter field around a compact binary can be described as a classical wave characterized by an amplitude, phase and direction [^1] We assume that the background vector field is linearly polarized. However, this is not a necessary condition for the local vector field in the halo. For example, in [32, 21] the authors perform small-scale structure simulations with vector field dark matter with different polarizations for the background field. that can be considered to be constant within a certain characteristic volume during a characteristic time interval. We will refer to such space-time volume as a coherent patch. The vector field in different coherent patches can have very different amplitudes, phases and directions. Our model is based on numerical simulations. In Refs. [21, 24], the authors study the merger of idealized halos, usually called solitons, to characterize small-scale structure formation with vector field dark matter. As was shown in such spin-1 halo-scale simulations, the system evolves into an approximately spherical configuration, that has a core surrounded by a dark matter density distribution characterized by interference patterns. The characteristic size of the interference granules is given by the de Broglie wavelength of the particle $\lambda_{dB}$. This is similar to what has been shown for scalar field dark matter models (see for instance Refs. [23, 33, 34]). Due to the evolution of the interference patterns, at a given point in space the field loses coherence after a time that can be estimated as $t_{coh}\sim\lambda_{dB}/V_{0}$, where $V_{0}$ is the local velocity of the dark matter field. These characteristic scales ($\lambda_{dB}$ and $t_{coh}$) define the typical space-time size of the coherent patches.
@@
where $V_{0}$ is the velocity dispersion of the halo, which can be estimated from the Virial velocity, e.g. $V_{0}\sim 10^{-3}$ for the Milky Way. On the other hand, for the systems that can be probed by LISA with a high signal-to-noise ratio, such as supermassive black holes of $M\sim 10^{6}\rm{M}_{\odot}$, the largest separations correspond [^2] In principle, LISA can detect binaries with lower masses at much larger separations (which would merge in the band of ground-based detectors), but at much lower signal-to-noise ratios, since the latter scales with the total mass. to the lowest frequencies LISA is sensitive to, $f\sim 10^{-5}\rm{Hz}$. Using Kepler’s third law we can calculate the semi-major axis of such systems, giving $r_{12}\sim 10^{8}\,\rm{km}$; and for the corresponding emitted gravitational wave $\lambda_{GW}\sim 10^{10}\,\rm{km}$. Furthermore, for $m_{A}<10^{-16}\rm{eV}$, we have that $t_{coh}>1\rm{yrs}$, so for the field masses considered in this work the patch remains coherent during a typical LISA observation, which ranges from 1 yrs to 4 yrs. In summary, we have that the de-Broglie wavelength is much larger than the orbital separation of the binary, and during the observation, the patch remains coherent. We therefore expect the background vector field (and its associated metric perturbation) to be approximately homogeneous at scales comparable to the size of the binary system, i.e. $\bar{A}^{\mu}$ and $h^{A}_{\mu\nu}$ are approximately independent of spatial coordinates and only depend on time. In Appendix A we give a more detailed discussion of the previous assumption, based on numerical simulations.
@@
where $\bar{g}_{\mu\nu}$ and $\bar{A}_{\mu}$ are the background metric and vector field respectively, while the metric perturbations $h^{A}_{\mu\nu}$ and $h^{M}_{\mu\nu}$ are produced by the vector field and the binary system respectively. The quantity $\delta A_{\mu}$ is a small perturbation that we will neglect henceforth, but that in principle includes inhomogeneities and the backreaction of the binary on the vector field. We here ignore the cosmological expansion of the universe and assume the background metric is the flat Minkowski spacetime [^3] The typical time scales of a gravitational-wave detector ($\sim$ yrs) are much smaller than the Hubble time in the matter era, and the main correction due to cosmic expansion is a redshift in the wavelength of the propagating gravitational wave [35]., $\bar{g}_{\mu\nu}=\eta_{\mu\nu}$. Thus, when we ignore the presence of the vector field, the metric perturbation $h^{M}_{\mu\nu}$ is nothing but the post-Newtonian solution for an inspiraling binary, as reviewed for example in [8].
@@
With this in hand, we then focus on the metric perturbation induced by the ordinary matter sector. In the absence of the vector field, the solution for $h_{\mu\nu}^{M}$ can be obtained through standard post-Newtonian methods [8], as discussed earlier. We then need to understand the backreaction of the vector field $\bar{A}^{\mu}$ on $h_{\mu\nu}^{M}$. The dominant correction to this metric perturbation is that induced by the correction to the trajectories of the binary system, which, as we will see, is effectively a conservative force induced by $\bar{A}^{\mu}$ that modifies the binary’s binding energy. By the balance law, this modification then alters the evolution of the orbital frequency, and thus, the gravitational-wave phase. This is the main and largest effect we consider in this paper, and thus, we leave the full back-reaction of $h_{\mu\nu}^{A}$ on $h_{\mu\nu}^{M}$ and vice versa to future work (see Appendix A for a discussion).
@@ III Evolution of the background ultralight dark-matter vector field and its metric perturbation
By considering a homogeneous field in flat spacetime we can further simplify the stress-energy tensor given in Eq. (4). First, we note that $-\frac{1}{2}F^{\mu\nu}F_{\mu\nu}+m_{A}^{2}A^{\mu}A_{\mu}=-\dot{A}^{2}+m_{A}^{2}A^{2}$ and $F^{\mu\gamma}F_{\nu\gamma}=\dot{A}^{2}$, where the square means the spatial norm in flat spacetime. Then, the components of the energy-momentum tensor for the homogeneous field are
@@
$-2\ddot{\psi}\,{\delta^{i}}_{j}-\partial^{(i}\dot{B}_{j)}-\frac{1}{2}{\ddot{E}^{i}}_{\,\,j}=8\pi G\,\rho_{A}\cos(2m_{A}t+\gamma){\hat{X}^{i}}_{\,\,j}\,.$ (15)

### v2

@@ I Introduction
One distinctive feature of ultralight dark-matter candidates is that they induce oscillatory gravitational potentials. The frequency of the oscillations is determined essentially by the mass of the ultralight dark-matter particles. In cosmological contexts, the dark-matter field oscillations usually average out, but this need not be the case around astrophysical systems. In fact, previous work has shown that an oscillatory ultralight, dark-matter environment induces oscillatory behavior in the spacetime metric of astrophysical systems, for example impacting the orbital dynamics and evolution of binary pulsars [Blas:2019hxz, Blas:2016ddr, LopezNacir:2018epg]. Could such an ultralight, dark-matter environment also affect the gravitational waves emitted by compact binaries?
General relativity predicts that gravitational waves are excited by any system with a non-trivial, time-varying multipole-moment structure, such as due to the orbital acceleration of a binary system. When in isolation, the orbital dynamics and evolution of inspiraling compact binaries can be determined from perturbative solutions to the Einstein equations. In practice, when the compact binary is sufficiently well-separated, one employs a post-Minkowskian expansion (in weak fields) or a post-Newtonian expansion (in both weak fields and small velocities) to calculate the gravitational waves in the inspiral phase of coalescence from first principles [Blanchet:2013haa]. When not in isolation, for example due to the presence of an accretion disk [Yunes:2011ws, Kocsis:2011dr, Speri:2022upm] or a dark-matter environment [Cole:2022yzw, Barausse:2014tra], the orbital dynamics and evolution of inspiraling compact binaries is in general modified, leaving an environmental imprint on the gravitational waves generated. The question is then not whether an ultralight, dark-matter environmental imprint is present in gravitational waves, but rather whether this imprint is detectable.
One of the most studied phenomena connecting dark-matter fields and black holes is the superradiant instability, which leads to the formation of a bosonic condensate around spinning black holes [Baumann:2018vus, Boskovic:2024fga, Baryakhtar:2017ngi, Pani:2012vp]. Such studies lead to the following exclusion windows for the mass of the vector field at 68% confidence limit [Stott:2020gjj]: $6.2\times 10^{-15}{\rm eV}\leq m\leq 3.9\times 10^{-11}{\rm eV},\,2.8\times 10^{-22}{\rm eV}\leq m\leq 1.9\times 10^{-16}{\rm eV}$, which incorporate observations of the event GW190521 and the shadow of M87*. In this paper we focus on a different phenomena than the one studied in previous work. In our set-up, for the sake of simplicity, we consider black holes with negligible spins so that the formation of a condensate via supperradiance can be neglected.
@@
Recent efforts have also been made to search for the direct imprint of ultralight dark matter in gravitational waves. For example, in [Brax:2024yqh, Marriott-Best:2025sez], the dephasing of gravitational waves produced by dark-matter fields was considered. In these works, the dephasing is produced in the propagation of the gravitational-wave signal from the source to the detector, caused by the gravitational potentials generated by the ultralight dark-matter fields. In this work, however, we focus on how these gravitational potentials create a dephasing in the generation of the gravitational-wave signal, by changing the world-line of the binary system components, rather than by correcting the way the gravitational waves propagate.
@@
When the binary systems emitting gravitational waves is immersed in a virialized ultralight, dark-matter halo, the dark matter that forms the halo (be it scalar or vector particles) can be described as a collection of waves with random phases [Amin:2022pzv, Centers:2019dyn, Schive:2014dra, Lopez-Sanchez:2025osk]. These waves induce oscillations in the spacetime metric, which then perturb the orbital motion of the binary. In particular, the ultralight dark-matter oscillations induce a correction to the acceleration of the binary components, which then affects the orbital energy and angular momentum, introducing dark-matter corrections to Kepler’s third law. The ultralight dark-matter environment, therefore, perturbs the orbital motion of the binary, which then leaves an imprint on the time-varying multipole-moment structure of the spacetime, and thus, on the gravitational waves emitted. Given this, one expects that the strength of the dark-matter imprint on the emitted gravitational waves to be proportional to the dark-matter density and sensitive to the dark-matter mass, which sets the dark-matter oscillation scale, specially close to resonances between this oscillation scale and the orbital timescale.
@@
As we show in the paper, the strength of the ultralight dark-matter effect on the gravitational waveform is proportional to the energy density $\rho_{\rm{DM}}$ of the dark-matter field in the surroundings of the black holes binary. Several works [local_dm_density_1, local_dm_density_2] have estimated the dark-matter density in our Solar System to be $\rho_{\rm{DM}}\sim 0.3\rm{GeV}/\rm{cm}^{3}\sim 0.01\rm{M}_{\odot}/\rm{pc}^{3}$. However, the dark-matter density distribution in galaxies is still a matter of debate. In particular, the dark-matter distribution surrounding black holes remains uncertain and varies depending on the dark-matter model. For example, various works [Alcubierre:2024mtq, Davies_2020, Cruz_Osorio_2011, Bar_2019, Aghaie:2023lan, Hancock:2025ois] suggest that scalar and vector field models in the particle regime ($\lambda_{c}\sim m^{-1}\ll r_{s}$, with $r_{s}$ the black hole Schwarzschild radius) spike near black holes, where the density may be several orders of magnitude bigger than the one estimated for our Solar System. In our work, we consider fields and black holes that are in the wave regime ($\lambda_{c}\gg r_{s}$), where the existence of a local density enhancement is unclear. We find that the vector ultralight dark-matter corrections are indeed proportional to the dark-matter density, and inversely proportional to a certain power of the ultralight dark-matter mass and the total mass of the binary. Moreover, we find that when the gravitational-wave frequency exceeds a certain threshold, given by a stationary point condition that depends on the vector field mass, the ultralight dark-matter corrections are strongly suppressed. On the other hand, when the gravitational-wave frequency is below this threshold, the vector ultralight dark-matter effect on the gravitational-wave phase for quasi-circular orbits is partially degenerate with the time and phase of coalescence of the binary, rendering this effect hard to measure. This leads to a Heaviside-type modification to the Fourier transform of the gravitational-wave phase, and it is this sudden turn-on/turn-off nature of the modification that yields the dark-matter effect potentially measurable.
Our work suggests that chirping, massive binaries that sample a range of frequencies in the LISA band (and thus, that cross the stationary point condition activating the Heaviside dark-matter modificatons) are necessary to observe this type of dark-matter effect. Through a Fisher analysis, we estimate that LISA would be able to constraint the ultralight dark-matter mass range $(10^{-19},10^{-16}){\rm{eV}}$ for gravitational waves emitted by massive $(10^{2},10^{6})\rm{M}_{\odot}$] black hole binaries of comparable mass, provided the ultralight dark-matter density is above $(10^{14},10^{15})\;\rm{M}_{\odot}\;{\rm{pc}}^{-3}$. Our work, therefore, suggests that future LISA observations of the gravitational waves emitted in the quasi-circular inspiral of massive black holes may detect or yield interesting constraints on the density of vector-type ultralight dark matter.
@@ II Ultralight dark-matter vector field around a compact binary
In this section, we present the vector-field dark-matter model that we study in this paper, following mostly [LopezNacir:2018epg], for completeness and to establish notation. Let us then describe the dark matter through an ultralight vector field minimally coupled to gravity, whose action is
$\mathcal{S}=\int dx^{4}\sqrt{-g}\left[R-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}-\frac{m_{A}^{2}}{2}A^{\mu}A_{\mu}+\mathcal{L}_{M}\right],$ (1)
@@
We consider that the vector dark matter field around a compact binary can be described as a classical wave characterized by an amplitude, phase and direction [^1] We assume that the background vector field is linearly polarized. However, this is not a necessary condition for the local vector field in the halo. For example, in [Amaral:2024tjg, Amin:2022pzv] the authors perform small-scale structure simulations with vector field dark matter with different polarizations for the background field. that can be considered to be constant within a certain characteristic volume during a characteristic time interval. We will refer to such space-time volume as a coherent patch. The vector field in different coherent patches can have very different amplitudes, phases and directions. Our model is based on numerical simulations. In Refs. [Amin:2022pzv, Lopez-Sanchez:2025osk], the authors study the merger of idealized halos, usually called solitons, to characterize small-scale structure formation with vector field dark matter. As was shown in such spin-1 halo-scale simulations, the system evolves into an approximately spherical configuration, that has a core surrounded by a dark matter density distribution characterized by interference patterns. The characteristic size of the interference granules is given by the de Broglie wavelength of the particle $\lambda_{dB}$. This is similar to what has been shown for scalar field dark matter models (see for instance Refs. [Schive:2014dra, Veltmaat:2018dfz, Liu:2022rss]). Due to the evolution of the interference patterns, at a given point in space the field loses coherence after a time that can be estimated as $t_{coh}\sim\lambda_{dB}/V_{0}$, where $V_{0}$ is the local velocity of the dark matter field. These characteristic scales ($\lambda_{dB}$ and $t_{coh}$) define the typical space-time size of the coherent patches.
@@
where $V_{0}$ is the velocity dispersion of the halo, which can be estimated from the Virial velocity, e.g. $V_{0}\sim 10^{-3}$ for the Milky Way. On the other hand, for the systems that can be probed by LISA with a high signal-to-noise ratio, such as supermassive black holes of $M\sim 10^{6}\rm{M}_{\odot}$, the largest separations correspond [^2] In principle, LISA can detect binaries with lower masses at much larger separations (which would merge in the band of ground-based detectors), but at much lower signal-to-noise ratios, since the latter scales with the total mass. to the lowest frequencies LISA is sensitive to, $f\sim 10^{-5}\rm{Hz}$. Using Kepler’s third law we can calculate the semi-major axis of such systems, giving
$r_{12}\sim 3.2\times 10^{8}\left(\frac{M}{10^{6}\rm M_{\odot}}\right)^{\frac{1}{3}}\left(\frac{f}{10^{-5}\rm Hz}\right)^{-\frac{2}{3}}\rm{km}\,$, (6)
and for the corresponding emitted gravitational wave $\lambda_{GW}\sim 10^{10}\,\rm{km}$. Furthermore, for $m_{A}<10^{-16}\rm{eV}$, we have that
$t_{coh}\sim 65\,\rm{yrs}\left(\frac{10^{-3}}{V_{0}}\right)^{2}\left(\frac{10^{-18}\rm{eV}}{m_{A}}\right)\,$, (7)
so for the field masses considered in this work the patch remains coherent during a typical LISA observation, which ranges from 1 yrs to 4 yrs. In summary, we have that the de-Broglie wavelength is much larger than the orbital separation of the binary, and during the observation, the patch remains coherent. We therefore expect the background vector field (and its associated metric perturbation) to be approximately homogeneous at scales comparable to the size of the binary system, i.e. $\bar{A}^{\mu}$ and $h^{A}_{\mu\nu}$ are approximately independent of spatial coordinates and only depend on time. In Appendix A we give a more detailed discussion of the previous assumption, based on numerical simulations.
@@
where $\bar{g}_{\mu\nu}$ and $\bar{A}_{\mu}$ are the background metric and vector field respectively, while the metric perturbations $h^{A}_{\mu\nu}$ and $h^{M}_{\mu\nu}$ are produced by the vector field and the binary system respectively. The quantity $\delta A_{\mu}$ is a small perturbation that we will neglect henceforth, but that in principle includes inhomogeneities and the backreaction of the binary on the vector field. We here ignore the cosmological expansion of the universe and assume the background metric is the flat Minkowski spacetime [^3] The typical time scales of a gravitational-wave detector ($\sim$ yrs) are much smaller than the Hubble time in the matter era, and the main correction due to cosmic expansion is a redshift in the wavelength of the propagating gravitational wave [Bieri:2017vni]., $\bar{g}_{\mu\nu}=\eta_{\mu\nu}$. Thus, when we ignore the presence of the vector field, the metric perturbation $h^{M}_{\mu\nu}$ is nothing but the post-Newtonian solution for an inspiraling binary, as reviewed for example in [Blanchet:2013haa].
@@
With this in hand, we then focus on the metric perturbation induced by the ordinary matter sector. In the absence of the vector field, the solution for $h_{\mu\nu}^{M}$ can be obtained through standard post-Newtonian methods [Blanchet:2013haa], as discussed earlier. We then need to understand the backreaction of the vector field $\bar{A}^{\mu}$ on $h_{\mu\nu}^{M}$. The dominant correction to this metric perturbation is that induced by the correction to the trajectories of the binary system, which, as we will see, is effectively a conservative force induced by $\bar{A}^{\mu}$ that modifies the binary’s binding energy. By the balance law, this modification then alters the evolution of the orbital frequency, and thus, the gravitational-wave phase. This is the main and largest effect we consider in this paper, and thus, we leave the full back-reaction of $h_{\mu\nu}^{A}$ on $h_{\mu\nu}^{M}$ and vice versa to future work (see Appendix A for a discussion).
@@ III Evolution of the background ultralight dark-matter vector field and its metric perturbation
By considering a homogeneous field in flat spacetime we can further simplify the stress-energy tensor given in Eq. (4). First, we note that $\frac{1}{2}F^{\mu\nu}F_{\mu\nu}+m_{A}^{2}A^{\mu}A_{\mu}=-\dot{A}^{2}+m_{A}^{2}A^{2}$ and $F^{\mu\gamma}F_{\nu\gamma}=\dot{A}^{2}$, where the square means the spatial norm in flat spacetime. Then, the components of the energy-momentum tensor for the homogeneous field are
@@
$2\ddot{\psi}\,{\delta^{i}}_{j}-\partial^{(i}\dot{B}_{j)}-\frac{1}{2}{\ddot{E}^{i}}_{\,\,j}=8\pi G\,\rho_{A}\cos(2m_{A}t+\gamma){\hat{X}^{i}}_{\,\,j}\,.$ (17)